# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 3. A splitting of symplectic K-theory
In this section, we record a motivic analogue of a classical splitting result. Let $\text{bsp}\simeq\tau_{\geq 0}\Sigma^{4}\text{KO}$ denote the connective cover of $\Sigma^{4}\text{KO}$. Then there is an equivalence of spectra [MM76, Theorem 1.5]:
@@ 3.1. Very effective slice filtration
The very effective slice filtration was introduced by Spitzweck–Østvær in [SØ12, Section 5] and further studied by Bachmann in [Bac17]. We follow the treatments given in [Ban25, Section 1] and [BH21b, Section 13].
@@
Lemma 3.2.
[Bac17, Lemma 8] For any $X\in\textup{SH}(F)$, we have
@@
$E_{2}=A_{*}(P_{n}Y)\implies A_{*}(Y)$
is by smashing the Whitehead tower for $Y$ with $A$ and applying homotopy groups. Now, let $E$ and $X$ be motivic spectra. Then there is a (generalized) very effective slice spectral sequence [CHR24]
$E_{2}=E_{**}(\tilde{\text{s}}_{n}X)\implies E_{**}(X)$
@@ 3.2. A splitting of symplectic K-theory
By definition, we have $\tilde{\text{f}}_{2}\text{KQ}\simeq\Sigma^{4,2}\text{ksp}$. Bachmann’s identification of the very effective slices of KQ [Bac17, Theorem 16] shows that
@@
Using the long exact sequence in homology and the isomorphisms from 2.10
@@
As we will see, this spectrum level splitting allows us to calculate $\pi_{**}^{F}\text{ksp}$ in a particularly nice way.
@@
There is another way to relate the motivic spectra ksp and kq that is relevant to our work. In [BH21a], Bachmann-Hopkins construct Adams operations $\psi^{q}$ on kq. In [KQ24], the homotopy groups of the spectrum L were computed, where L sits in a cofiber sequence
@@
However, Bachmann-Hopkins show that the map $\psi^{3}-1$ factors through the second very effective cover of kq. At least 2-locally, this defines a motivic image of J spectrum $\text{j}_{\text{O}}$, which sits in a cofiber sequence
@@
In future work, we will show that the homotopy groups $\pi_{**}^{F}\text{j}_{\text{O}}$ are isomorphic to the 0 and 1-lines of the kq-resolution.
@@ 4. The inductive process to calculate $\pi_{**}^{F}(\mathrm{kq}\otimes\mathrm{kq})$
In this section, we consider the motivic Adams spectral sequence converging to $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$. We show that the $E_{2}$-page splits using motivic Brown–Gitler comodules, then outline an inductive process to compute the $E_{2}$-page by a series of algebraic Atiyah–Hirzebruch spectral sequences.
@@ 4.1. The motivic Adams spectral sequence for $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$
[…]
$E_{2}^{s,f,w}=\text{Ext}^{s,f,w}_{\mathcal{A}_{F}^{\vee}}(\textup{H}_{**}(\textup{kq}\otimes\textup{kq}))\implies\pi_{s,w}^{F}(\textup{kq}\otimes\textup{kq}),\quad d_{r}:E_{r}^{s,f,w}\to E_{r}^{s-1,f+r,w}.$
We can rewrite the $E_{2}$-page of this spectral sequence in a convenient way.
@@
[…]
$E_{2}^{s,f,w}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(\Sigma^{4k,2k}B_{0}^{F}(k)).$
@@
By 2.10 and Proposition 2.12, we have an isomorphism of $\mathcal{A}^{\vee}_{F}$-comodule algebras
@@
[…]
$E_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{F}}((\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))^{\vee}_{F}).$
Finally, using 2.19 and the fact that $\mathbb{M}_{2}^{F}$ is compact as an $\mathcal{A}(1)^{\vee}_{F}$-comodule, we can rewrite this as
$E_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{F}(k)),$
@@
Up until this point, all computation has been a precise motivic analogue of the classical work of Mahowald [Mah81]. However, Mahowald’s computation of $\pi_{*}(\text{bo}\otimes\text{bo})$ heavily relies on the fact that the classical integral Brown–Gitler spectra $\textup{H}\mathbb{Z}_{k}$ give rise to a topological splitting:
@@
4.2. The algebraic Atiyah–Hirzebruch spectral sequence
We now wish to compute the groups $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(k))$ appearing in the decomposition of the $E_{2}$-page from 4.1. To do this, one can use a series of algebraic Atiyah–Hirzebruch spectral sequences (aAHSS) along with the short exact sequences of Brown–Gitler comodules from Proposition 2.21. This is originally inspired by the work of Behrens-Ormsby-Stapleton-Stojanoska [BOSS19]. We explain the method here for clarity:
@@ 4.2. The algebraic Atiyah–Hirzebruch spectral sequence
    $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})\otimes\mathbb{M}_{2}^{F}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1));$
  2. (2)
@@
Remark 4.3.
Since we are only concerned with using the kq-resolution to compute the $v_{1}$-periodic component of $\pi_{**}^{F}\mathbb{S}$, passing to $v_{1}$-torsion free Ext groups loses no information. Additionally, we will see that there are multiple $v_{1}$-torsion classes arising in these spectral sequences, and ignoring these classes drastically simplifies our calculations.
[…]
$E_{2}=\text{Ext}_{\mathcal{A}^{\vee}_{F}}(\textup{H}_{**}(\textup{kq}))\cong\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F}).$
There are explicit formulae for this algebra in $\mathbb{C}$-motivic [IS11, Theorem 6.6], $\mathbb{R}$-motivic [Hil11, Theorem 4.8], [GHIR20, Theorem 6.2], and $\mathbb{F}_{q}$-motivic homotopy [Kyl15, Theorems 4.4.2, 4.4.3]. We can use $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})$ as input for an algebraic Atiyah–Hirzebruch spectral sequence $\textbf{aAHSS}(B_{0}^{F}(1))$ computing the next summand, $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1))$, which we now outline.
00{0}${{\text{F}_{0}B_{0}^{F}(1)}}$${{\text{F}_{1}B_{0}^{F}(1)}}$${{\text{F}_{2}B_{0}^{F}(1)}}$${{\text{F}_{3}B_{0}^{F}(1)=B_{0}^{F}(1)}}$${{\mathbb{M}_{2}^{F}\{[1]\}}}$00{0}${{\mathbb{M}_{2}^{F}\{[\overline{\xi}_{1}]\}}}$${{\mathbb{M}_{2}^{F}\{[\overline{\tau}_{1}]\}}}$
We refer to this filtration as the cellular filtration. The $\textbf{aAHSS}(B_{0}^{F}(1))$ comes from the exact couple arising from applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(-)$ to this filtration. This spectral sequence has signature
$E_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})\otimes\mathbb{M}^{F}_{2}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(1)).$
The $E_{1}$-page of this spectral sequence consists of 3 copies of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ in Atiyah–Hirzebruch filtration degrees 0, 2, and 3, which we index by $a$. Each Atiyah–Hirzebruch filtration piece is suitably shifted by the degree of the cell it is attached to. To make this precise, we can rewrite the $E_{1}$-page as
$E_{1}\cong\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})\oplus\Sigma^{2,1}\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})\oplus\Sigma^{3,1}\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F}).$
We give a heuristic for this spectral sequence in Fig. 2.
$[1]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})$$[\overline{\xi}_{1}]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})$$[\overline{\tau}_{1}]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})$
Figure 2. A heuristic for the aAHSS$(B_{0}^{F}(1))$.
If $x\in\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ has degree $(s,f,w,a)$, then $|\Sigma^{p,q}x|=(s+p,f,w+q,a).$ We will let $\alpha[i]$ denote the copy of the class $\alpha\in\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})$ in Atiyah–Hirzebruch filtration $i$. With this notation, differentials in the $\textbf{aAHSS}(B_{0}^{F}(1))$ are of the form
$d_{r}:E_{r}^{s,f,w,a}\to E_{r}^{s-1,f+1,w,a-r}.$
@@
The differentials in this spectral sequence are computed by the cobar complex: to determine the differential on a class $\alpha[k]$, we first lift to a representative in the cobar complex for $\text{F}_{k}B_{0}^{F}(1)$, compute the cobar differential on this representative, then project down to the appropriate filtration. In other words, the differentials witness the attaching maps in the algebraic cell complex for $B_{0}^{F}(1)$. The $d_{1}$-differential witnesses the $h_{0}$-attaching map between the cells $[\overline{\tau}_{1}]$ and $[\overline{\xi}_{1}]$, so is nonzero only on Atiyah–Hirzebruch filtration 3 and lands in Atiyah–Hirzebruch filtration 2. This gives the following formula:
@@
Linearity over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ combined with the above formula gives the behavior of $d_{1}$. The $d_{2}$-differential witness the $h_{1}$-attaching map between the cells $[\overline{\xi}_{1}]$ and $[1]$, so is nonzero only on Atiyah–Hirzebruch filtration 2 and lands in Atiyah–Hirzebruch filtration 0. This gives the following formula:
@@
We will see that this differential is dependent on the base field $F$, as is evident from the formula. By inspection of the cellular filtration on $B_{0}^{F}(1)$, we see the following.
@@
In the $\textup{{aAHSS}}(B_{0}^{F}(1))$, we have that $E_{4}=E_{\infty}$.
The process of solving extension problems on the $E_{\infty}$-page is more involved, so we postpone the discussion until we specialize to particular fields $F$.
@@
Though not tantamount to computing the ring of cooperations, it still provides insight to compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1)^{\otimes i})$. Classically, Mahowald is able to express this group in terms of Adams covers of bo and bsp. This is not quite the case here; see Proposition 5.5 and 7.10.
Now, we discuss the second step of our method. While one could analyze the algebraic Atiyah–Hirzebruch spectral sequence arising from the cellular filtration on $B_{0}^{F}(k)$, this spectral sequence becomes drastically more complicated. We will proceed in a different manner. After finding a formula for $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(1))$, one can induct along the short exact sequences of Proposition 2.21 to compute $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(k))$. If $k$ is even, then there is a short exact sequence
$0\to\Sigma^{2k,k}B_{0}^{F}(\tfrac{k}{2})\to B_{0}^{F}(k)\to B_{1}^{F}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0$
Applying $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(-)$, we see that modulo $v_{1}$-torsion the connecting homomorphism is trivial. This reduces the computation of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(k))$ to the kernel and cokernel in these short exact sequences in Ext. The kernels are taken care of by the inductive hypothesis, and the cokernels are a simple computation involving a change of rings isomorphism. This finishes the computation when $k$ is even.

### v2

@@ 3. A splitting of symplectic K-theory
In this section, we record a motivic analogue of a classical splitting result. Let $\text{bsp}\simeq\tau_{\geq 0}\Sigma^{4}\text{KO}$ denote the connective cover of $\Sigma^{4}\text{KO}$. Then there is an equivalence of spectra [40, Theorem 1.5]:
@@ 3.1. Very effective slice filtration
The very effective slice filtration was introduced by Spitzweck–Østvær in [53, Section 5] and further studied by Bachmann in [6]. We follow the treatments given in [8, Section 1] and [4, Section 13].
@@
Lemma 3.1 ([6, Lemma 8]).
For any $X\in\textup{SH}(F)$, we have
@@
$\textup{E}_{2}=A_{*}(P_{n}Y)\implies A_{*}(Y)$
is by smashing the Whitehead tower for $Y$ with $A$ and applying homotopy groups. Now, let $E$ and $X$ be motivic spectra. Then there is a (generalized) very effective slice spectral sequence [16]
$\textup{E}_{2}=E_{**}(\tilde{\text{s}}_{n}X)\implies E_{**}(X)$
@@ 3.2. A splitting of symplectic K-theory
By definition, we have $\tilde{\text{f}}_{2}\text{KQ}\simeq\Sigma^{4,2}\text{ksp}$. Bachmann’s identification of the very effective slices of KQ [6, Theorem 16] shows that
@@
Using the long exact sequence in homology and the isomorphisms from 2.5
@@
As we will see, this spectrum level splitting allows us to calculate $\pi_{**}^{F}(\text{ksp})$ in a particularly nice way.
@@
There is another way to relate the motivic spectra ksp and kq that is relevant to our work. In [3], Bachmann–Hopkins construct Adams operations $\psi^{q}$ on kq. In [32], the homotopy groups of the spectrum L were computed, where L sits in a cofiber sequence
@@
However, Bachmann–Hopkins show that the map $\psi^{3}-1$ factors through the second very effective cover of kq. At least 2-locally, this defines a motivic image of J spectrum $\text{j}_{\text{O}}$, which sits in a cofiber sequence
@@
Culver–Quigley show that the $\mathbb{C}$-motivic 0- and 1-lines of the kq-resolution are isomorphic to $\pi_{**}^{\mathbb{C}}(\text{j}_{\text{O}})$ [17, Theorem 7.13]. We imagine that the same is true over arbitrary base fields.
@@ 4. The inductive process to calculate $\pi_{**}^{F}(\mathrm{kq}\otimes\mathrm{kq})$
In this section, we consider the motivic Adams spectral sequence converging to $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$. We show that the $\textup{E}_{2}$-page splits using motivic Brown–Gitler comodules, then outline an inductive process to compute the $\textup{E}_{2}$-page by a series of algebraic Atiyah–Hirzebruch spectral sequences.
@@ 4.1. The motivic Adams spectral sequence for $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$
There is an $\textup{H}\mathbb{F}_{2}$-motivic Adams spectral sequence computing the cooperations (mASSF($\textup{kq}\otimes\textup{kq}$)). Since $\text{H}\mathbb{F}_{2}$ is flat in the sense of Adams, this takes the form [28]:
$\textup{E}_{2}^{s,f,w}=\text{Ext}^{s,f,w}_{\mathcal{A}_{F}^{\vee}}(\textup{H}_{**}(\textup{kq}\otimes\textup{kq}))\implies\pi_{s,w}^{F}(\textup{kq}\otimes\textup{kq}),\quad d_{r}:\text{E}_{r}^{s,f,w}\to\text{E}_{r}^{s-1,f+r,w}.$
We can rewrite the $\textup{E}_{2}$-page of this spectral sequence in a convenient way.
@@
The $\textup{E}_{2}$-page of the mASSF($\textup{kq}\otimes\textup{kq}$) is given by
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(\Sigma^{4k,2k}B_{0}^{F}(k)).$
@@
By 2.5 and Proposition 2.7, we have an isomorphism of $\mathcal{A}^{\vee}_{F}$-comodule algebras
@@
Combined with a standard change of rings argument [49, Appendix A], we can rewrite the $\textup{E}_{2}$-page of the mASSF($\textup{kq}\otimes\textup{kq}$) as
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{F}}((\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))^{\vee}_{F}).$
Finally, using 2.9 and the fact that $\mathbb{M}_{2}^{F}$ is compact as an $\mathcal{A}(1)^{\vee}_{F}$-comodule, we can rewrite this as
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{F}(k)),$
@@
Up until this point, all computation has been a precise motivic analogue of the classical work of Mahowald [41]. However, Mahowald’s computation of $\pi_{*}(\text{bo}\otimes\text{bo})$ heavily relies on the fact that the classical integral Brown–Gitler spectra $\textup{H}\mathbb{Z}_{k}$ give rise to a topological splitting:
@@
4.2. Computing the summands of the $\textbf{mASS}^{F}(\text{kq}\otimes\text{kq})$
We now wish to compute the groups $\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(k))$ appearing in the decomposition of the $\textup{E}_{2}$-page from 4.1. To do this, one can use a series of algebraic Atiyah–Hirzebruch spectral sequences (aAHSS) along with the short exact sequences of Brown–Gitler comodules from Proposition 2.11. This is originally inspired by the work of Behrens–Ormsby–Stapleton–Stojanoska [10]. We explain the method here for clarity:
@@ 4.2. Computing the summands of the $\textbf{mASS}^{F}(\text{kq}\otimes\text{kq})$
    Compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B^{F}_{0}(1))/(v_{1}\text{-torsion})$ using an aAHSS:
    $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})\otimes\mathbb{M}_{2}^{F}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1));$
  2. (2)
    Compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B^{F}_{0}(k))/(v_{1}\text{-torsion})$ by induction and Proposition 2.11.
Remark 4.2.
Since we are only concerned with using the kq-resolution to compute the $v_{1}$-periodic component of $\pi_{**}^{F}(\mathbb{S})$, passing to $v_{1}$-torsion free Ext groups loses no information. Additionally, we will see that there are multiple $v_{1}$-torsion classes arising in these spectral sequences, and ignoring these classes drastically simplifies our calculations.
4.2.1. The algebraic Atiyah–Hirzebruch spectral sequence
We have that $B_{0}^{F}(0)\cong\mathbb{M}_{2}^{F}$, and so the first summand of the $\textbf{mASS}^{F}(\textup{kq}\otimes\textup{kq})$ is given by $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$. Notice that since $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{F}^{\vee}$ by 2.5, a standard change of rings argument shows that the $\textup{E}_{2}$-page of the $\textbf{mASS}^{F}(\textup{kq})$ is given by
$\textup{E}_{2}=\text{Ext}^{***}_{\mathcal{A}^{\vee}_{F}}(\textup{H}_{**}(\textup{kq}))\cong\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F}).$
There are explicit formulae for this algebra in $\mathbb{C}$-motivic [29, Theorem 6.6], $\mathbb{R}$-motivic [26, Theorem 4.8], [24, Theorem 6.2], and $\mathbb{F}_{q}$-motivic homotopy [34, Theorems 4.4.2, 4.4.3]. We can use $\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})$ as input for an algebraic Atiyah–Hirzebruch spectral sequence $\textbf{aAHSS}(B_{0}^{F}(1))$ computing the next summand, $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1))$, which we now outline.
@@ 4.2.1. The algebraic Atiyah–Hirzebruch spectral sequence
We refer to this filtration as the Atiyah–Hirzebruch filtration. The $\textbf{aAHSS}(B_{0}^{F}(1))$ comes from the exact couple arising from applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(-)$ to (4.3). This spectral sequence has signature
$\textup{E}_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})\otimes\mathbb{M}^{F}_{2}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(1)).$
The $\textup{E}_{1}$-page of this spectral sequence consists of 3 copies of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ in Atiyah–Hirzebruch filtration degrees 0, 2, and 3, which we index by $a$. Each Atiyah–Hirzebruch filtration piece is suitably shifted by the degree of the cell it is attached to. To make this precise, we can rewrite the $\textup{E}_{1}$-page as
$\textup{E}_{1}\cong\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})\{[1]\}\oplus\Sigma^{2,1}\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})\{[\overline{\xi}_{1}]\}\oplus\Sigma^{3,1}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})\{[\overline{\tau}_{1}]\}.$
We give a heuristic for this spectral sequence in Figure 4.1.
$1]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})[$$\overline{\xi}_{1}]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})[$$\overline{\tau}_{1}]\otimes\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{M}_{2})$
@@
If $x\in\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ has aAHSS degree $(s,f,w,a)$, then $|\Sigma^{p,q}x|=(s+p,f,w+q,a).$ We will let $\alpha[i]$ denote the copy of the class $\alpha\in\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(\mathbb{M}_{2}^{F})$ in Atiyah–Hirzebruch filtration $i$. With this notation, differentials in the $\textbf{aAHSS}(B_{0}^{F}(1))$ are of the form
$d_{r}:\text{E}_{r}^{s,f,w,a}\to\text{E}_{r}^{s-1,f+1,w,a-r}.$
@@
The differentials in this spectral sequence are computed by the cobar complex: to determine the differential on a class $\alpha[k]$, we first lift to a representative in the cobar complex for $\text{F}_{k}B_{0}^{F}(1)$, compute the cobar differential on this representative, then project down to the appropriate filtration. In other words, the differentials witness the attaching maps in the algebraic cell complex for $B_{0}^{F}(1)$.
The $d_{1}$-differential witnesses the $h_{0}$-attaching map between the cells $\overline{\tau}_{1}]$ and $\overline{\xi}_{1}]$, so is nonzero only on Atiyah–Hirzebruch filtration 3 and lands in Atiyah–Hirzebruch filtration 2. This gives the following formula:
@@
Linearity over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$ combined with the above formula gives the behavior of $d_{1}$. The $d_{2}$-differential witness the $h_{1}$-attaching map between the cells $\overline{\xi}_{1}]$ and $[1]$, so is nonzero only on Atiyah–Hirzebruch filtration 2 and lands in Atiyah–Hirzebruch filtration 0. This gives the following formula:
@@
We will see that this differential is dependent on the base field $F$, as is evident from the formula.
By inspection of the length of (4.3), we see the following.
@@
In the $\textup{{aAHSS}}(B_{0}^{F}(1))$, we have that $\textup{E}_{4}=\textup{E}_{\infty}$.
The process of solving extension problems on the $\textup{E}_{\infty}$-page is more involved, so we postpone the discussion until we specialize to particular fields $F$.
@@
Though not tantamount to computing the ring of cooperations, it still provides insight to compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(1)^{\otimes i})$. Classically, Mahowald is able to express this group in terms of Adams covers of bo and bsp. This is not quite the case here; see Proposition 5.4 and 7.6.
4.2.2. Ext of higher integral Brown–Gitler comodules
Now, we discuss the second step of our method. While one could analyze the algebraic Atiyah–Hirzebruch filtration on $B_{0}^{F}(k)$, the resulting spectral sequence becomes drastically more complicated. We will proceed in a different manner. After finding a formula for $\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(1))$, one can induct along the short exact sequences of Proposition 2.11 to compute $\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}(B_{0}^{F}(k))$.
When $k$ is even, then there is a short exact sequence
$0\to\Sigma^{2k,k}B_{0}^{F}(\tfrac{k}{2})\to B_{0}^{F}(k)\to B_{1}^{F}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0.$
Applying $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(-)$ gives a long exact sequence of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(\mathbb{M}_{2}^{F})$-modules. By a change of rings isomorphism, the Ext group of the cokernel may be rewritten:
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B_{1}^{F}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{F})\cong\text{Ext}_{\mathcal{A}(0)^{\vee}_{F}}^{***}(B_{1}^{F}(\tfrac{k}{2}-1)).$
Since $v_{1}$ acts trivially over $\mathcal{A}(0)^{\vee}$, we see that, modulo $v_{1}$-torsion, the connecting homomorphism is trivial. This reduces the computation of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(k))$ to the kernel and cokernel in these short exact sequences in Ext. The kernels are taken care of by the inductive hypothesis, and the cokernels are much simpler after reducing to $\mathcal{A}(0)^{\vee}_{F}$. This finishes the computation when $k$ is even.